A mother is five times as old as her son If twice the age of the son is 3 years less than half of the age of the mother how old are they
Answer by mananth(16946)

let son be x years
mother will be 5x years.
...
2x = (1/2)*5x-3
2x=5x/2 -3
2x=(5x-6)/2
4x=5x-6
5x-4x=6
x= 6 son's age
mother is 5*6= 30 years
...
m.ananth@hotmail.ca